1.00 mL of 12.0 M HCl is added to 1.00 L of a buffer that contains  0.110 M HNO2 and 0.150 M NaNO2. How many moles of HNO2 and NaNO2remain in solution after addition of the HCl?


1
Expert's answer
2021-10-01T10:59:02-0400

Number of moles of NaNO2 = "Volume \;of \;solution(in \;L) \times \;Molarity \;of\; NaNO2 = 1L \times 0.150 = 0.150 \;moles"


Number of moles of HNO2 = "Volume\; of\; solution(in L) \times \;Molarity\; of \;HNO2 = 1L \times 0.110 = 0.110\; moles"


Number of moles of HCl added = "Volume \;of \;HCl \times Molarity\; of \;HCl = \frac{1}{1000} \times12M = 0.012 \;moles"


Added HCl will react with NaNO2 to form

HCl + NaNO2 ------ NaCl + HNO2


Moles of NaNO2 remained in solution = initial moles - number of moles of HCl = 0.150 - 0.012 = 0.138 moles


Moles of HNO2 remained in solution = initial moles + number of moles of HCl = 0.110 + 0.012 = 0.122 moles